OpenStudy (anonymous):

Write the equation in slope-intercept and standard form for the line that goes through the points (1,3) and (2,5). What is the slope of the equation that is perpendicular to the line above?

OpenStudy (cwrw238):

do you know how to find the slope of this line?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

m=2. y=2x+1, then to find perpendicular we need the slopes to multiply to -1 (I used -1/2) so the equation of the perp. line would be y=-(1/2)x+1

OpenStudy (anonymous):

to find the slope (how I got m) is delta y over delta x. (5-3) / (2-1) = 2/1 = 2